# LPNSR: Optimal Noise-Guided Diffusion Image Super-Resolution Via Learnable Noise Prediction

### 12 Apr 2026 v4

@@ Abstract
Diffusion-based image super-resolution (SR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) observations, yet faces a fundamental trade-off between inference efficiency and reconstruction quality in limited-step sampling scenarios. A critical yet underexplored question is: what is the optimal noise to inject at each intermediate diffusion step? In this paper, we establish a theoretical framework that derives the closed-form analytical solution for optimal intermediate noise in diffusion models from a maximum likelihood estimation perspective, revealing a consistent conditional dependence structure that generalizes across diffusion paradigms. We instantiate this framework under the residual-shifting diffusion paradigm and accordingly design an LR-guided multi-input-aware noise predictor to replace random Gaussian noise. We further mitigate initialization bias with a high-quality pre-upsampling network. The compact 4-step trajectory uniquely enables end-to-end optimization of the entire reverse chain, which is computationally prohibitive for conventional long-trajectory diffusion models. Extensive experiments demonstrate that LPNSR achieves state-of-the-art perceptual performance on both synthetic and real-world datasets, without relying on any large-scale text-to-image priors. The source code of our method can be found at https://github.com/Faze-Hsw/LPNSR.
@@ 1 Introduction
We instantiate this framework on the residual-shifting diffusion paradigm, as its compact 4-step trajectory uniquely enables end-to-end optimization of the full reverse chain—computationally infeasible for long-trajectory models like DDPM(Ho et al., 2020). This allows us to learn a deep neural network for the theoretically optimal noise without modifying the pretrained denoising network or breaking the original efficient residual-shifting mechanism. We further address initialization bias from bicubic upsampling, a critical bottleneck for few-step sampling where limited iterations cannot correct initial deviations, via a pretrained pre-upsampling network to generate a high-quality initial state before diffusion. An additional key benefit of this design is that it enables arbitrary-step super-resolution inference without redesigning the diffusion framework’s hyper parameters or retraining the pretrained denoising network.
@@ 2 Related work
Image Super-Resolution. Along with the proliferation of deep learning, deep learning-driven approaches have progressively emerged as the dominant paradigm for SR(Dong et al., 2015; Rojas Sedó, 2022). Early prominent works primarily focused on training regression models using paired LR-HR data(Ahn et al., 2018; Kim et al., 2016; Wang et al., 2015). Though these models effectively capture the expectation of the posterior distribution, they inherently suffer from over-smoothing artifacts in generated results(Ledig et al., 2017; Menon et al., 2020; Sajjadi et al., 2017). To enhance the perceptual quality of reconstructed HR images, generative SR models have garnered growing interest—including autoregressive architectures(Dahl et al., 2017; Menick and Kalchbrenner, 2018; Van den Oord et al., 2016; Parmar et al., 2018). Despite notable gains in perceptual performance, autoregressive models typically incur substantial computational overhead. Additionally, GAN-based SR methods have attained remarkable success in perceptual quality(Guo et al., 2022; Karras et al., 2017; Ledig et al., 2017; Menon et al., 2020; Sajjadi et al., 2017), yet the training process remains unstable. More recently, diffusion-based models have become a focal point of SR research(Choi et al., 2021; Chung et al., 2022c; Kawar et al., 2022; Rombach et al., 2022; Saharia et al., 2023). These methods generally fall into two categories: those that concatenate the LR image to the denoiser’s input(Rombach et al., 2022; Saharia et al., 2023), and those that adapt the backward process of a pre-trained diffusion model(Choi et al., 2021; Chung et al., 2022c; Kawar et al., 2022). While these diffusion-based approaches yield promising performance, their methods still introduce unconstrained random Gaussian noise in each step of the reverse diffusion process, rather than meaningful noise maps.
@@
Diffusion Inversion. This paradigm aims to find the optimal noise maps that reconstruct the target image when fed into a diffusion model. Early works optimized text embeddings for better alignment(Gal et al., 2022; Mokady et al., 2023), and follow-up works further refined inversion via textual/visual prompts(Miyake et al., 2025; Nguyen et al., 2023) or intermediate noise map optimization(Ju et al., 2023; Kang et al., 2024; Meiri et al., 2023; Wallace et al., 2023). However, existing diffusion inversion methods are mostly heuristic step-wise optimization schemes tailored for image editing tasks. None of them establish a unified, generalizable theoretical derivation paradigm for optimal intermediate noise, nor provide a closed-form analytical solution that can generalize across mainstream diffusion paradigms. For SR tasks, InvSR(Yue et al., 2025) extended diffusion inversion to SR, but it is constrained by the long sampling trajectory of DDPM(Ho et al., 2020), which makes full-chain end-to-end optimization computationally prohibitive. It only optimizes the initial noise map, and fails to solve the fundamental problem of defining and optimizing the optimal intermediate noise for the full reverse sampling chain.
@@ 3.2 Theoretical derivation of optimal intermediate noise
Following standard diffusion posterior derivation practices(Ma et al., 2023), we adopt a non-informative prior $p_{\theta}\left(x_{0}|y\right)\propto 1$, and the marginal likelihood $p_{\theta}\left(x_{t-1}|y\right)$ is a normalization constant independent of $x_{0}$. This simplifies the posterior to $p_{\theta}\left(x_{0}|x_{t-1},y\right)\propto q\left(x_{t-1}|x_{0},y\right)$. Substituting the forward marginal Gaussian distribution in Eq. (2) and taking the logarithm, the log-likelihood becomes
$\log p_{\theta}\left(x_{0}|x_{t-1},y\right)\propto-\frac{1}{2\Sigma_{t-1}^{m}}\left\|x_{0}-\left(\boldsymbol{\mu}_{t-1}^{m}\right)^{-1}\left(x_{t-1}\right)\right\|^{2},$ (7)
where $\left(\boldsymbol{\mu}_{t-1}^{m}\right)^{-1}\left(\cdot\right)$ is the inverse mapping of the marginal mean function with respect to $x_{0}$. Maximizing the log-likelihood is equivalent to minimizing the above L2 norm, leading to the optimization objective:
$z_{t-1}^{*}=\arg\min_{z_{t-1}}\left\|x_{0}-\left(\boldsymbol{\mu}_{t-1}^{m}\right)^{-1}\left(x_{t-1}(z_{t-1})\right)\right\|^{2}.$ (8)
@@ 3.3 Instantiation to residual-shifting diffusion SR paradigm
We instantiate our general optimal noise derivation paradigm proposed in Sec. 3.2 to the residual-shifting diffusion frameworkYue et al. (2023), a efficient SR pipeline with a compact 4-step sampling trajectory. Unlike long-trajectory models (e.g., DDPMHo et al. (2020)) requiring hundreds of steps, this 4-step design uniquely enables end-to-end optimization of the full reverse chain, which is computationally prohibitive for conventional long-trajectory diffusion models.
@@ 4.1 Experimental setup
Training Details. We train the noise predictor on the LSDIRLi et al. (2023) dataset and the first 10k face images from the FFHQKarras et al. (2019) dataset for over 200k iterations, randomly cropping an image patch with a resolution of $256\times 256$ from the source image and synthesizing the LR image using the pipeline of RealESRGANWang et al. (2021) at each iteration. We adopt the AdamWLoshchilov and Hutter (2017) optimizer with a learning rate of $5\times 10^{-5}$ and a batch size of 16, while using the CosineAnnealingLoshchilov and Hutter (2016) scheduler with a minimum learning rate of $1\times 10^{-5}$. The hyperparameters for the loss function are set as $\lambda_{1}=1.0$, $\lambda_{l}=1.0$, and $\lambda_{g}=0.1$. During training, we set $T=4$ to remain consistent with ResShiftYue et al. (2023),and the noise variance hyperparameter $\kappa=2.0$ as well as the shifting sequence $\{\eta_{t}\}_{t=1}^{T}$ also follow the identical settings. The denoising network $f_{\theta}$ and the VQGAN autoencoderEsser et al. (2021) is frozen during training, only the noise predictor is optimized.
@@ 4.2 Experimental results
Inference Steps. Tab. 1 compares our PreSet-A (noise predictor only) and PreSet-B (with pre-upsampling) methods against original ResShiftYue et al. (2023) across 1-4 sampling steps. We observe that the pre-upsampling module delivers more significant gains with fewer steps, while our noise predictor stably improves perceptual performance across all settings, with all methods performing better with more steps. Regarding inference latency, our noise predictor introduces no noticeable overhead as its computation is performed entirely in the latent space. In contrast, the pre-upsampling network operates in the image space and thus incurs additional runtime. For 3-4 step inference, the model has sufficient iterations to mitigate initialization bias, so the pre-upsampling module can be optionally omitted to trade off a minor performance drop for faster inference. For 1-2 step inference, however, the pre-upsampling module is indispensable, as the limited sampling steps cannot compensate for initialization deviation without high-quality pre-upsampling. Qualitative results in Fig. 1 confirm that PreSet-B maintains consistent visual quality even in 1-2 step inference, while PreSet-A suffers from blurriness in low-step settings. We fix 4-step inference with pre-upsampling as our default LPNSR configuration for all following experiments.
@@
Performance Comparison. Tab. 2 and 3 presents a comprehensive comparison of our LPNSR against recent SOTA methods on the ImageNet-Test, RealSR and RealSet80 datasets. Compared to the baseline ResShiftYue et al. (2023), our LPNSR achieves remarkable improvements in perceptual metrics (e.g., NIQE, CLIPIQA, MUSIQ) while maintaining competitive fidelity. Against T2I-utilizing models such as OSEDiffWu et al. (2024a), InvSRYue et al. (2025) and SeeSRWu et al. (2024b), LPNSR delivers comparable or better perceptual quality without leveraging any pre-trained text-to-image priors. It also outperforms multi-step diffusion methods (e.g.,SeeSRWu et al. (2024b)) on core perceptual metrics. On real-world datasets, LPNSR ranks among the top-tier SOTA methods. It achieves leading perception-oriented metrics, such as NIQE, PI, CLIPIQA and MUSIQ on RealSR. On RealSet80, LPNSR attains top-2 PI score, the best MUSIQ and CLIPIQA scores among all competing methods. Qualitatively, Fig. 4 shows LPNSR generates sharper textures and more consistent structures than other methods, free from spurious details or over-smoothing (see Appendix for more visual comparisons). LPNSR generates SR images with sharp details, intact structural consistency, and no noticeable artifacts. It effectively restores natural textures and clear edge contours that align with the input LR structure, delivering visually coherent and realistic results.
@@
Analysis and Ablation Study of Noise Predictor. Across the 4-step coarse-to-fine denoising trajectory, our LR-guided noise predictor implements progressive prior guidance aligned with the denoising logic. As shown in Fig. 2, the predicted noise maps are highly aligned with the LR image’s structure and texture, following a hierarchical guidance pattern: the step 4 map anchors global structure to avoid initial sampling deviation, the step 3 map focuses on mid-frequency texture refinement to suppress cumulative error, and the step 2 map targets local fine-grained details to optimize perceptual quality. Statistical distribution analysis in Fig. 3 further validates this progressive, LR-aligned variation pattern, and confirms that the optimized noise distributions at all steps are not unconstrained random Gaussian distributions, verifying the intrinsic working mechanism of our predictor. We further conduct a step-wise ablation study on the RealSR dataset (see Tab. 4) to quantitatively verify the independent contribution of each step’s predictor: the full LPNSR model achieves the best overall performance, validating the effectiveness of our full-stage prior guidance; removing the step 4 predictor degrades both fidelity and perceptual quality, disabling the step 3 predictor causes the most severe PSNR drop, and replacing the step 2 predictor with random noise sharply impairs perceptual metrics, fully consistent with our qualitative and statistical observations.

### 14 Apr 2026 v5

@@ Abstract
Diffusion-based image super-resolution (SR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) observations. However, the inherent randomness injected during the reverse diffusion process causes the performance of diffusion-based SR models to vary significantly across different sampling runs, particularly when the sampling trajectory is compressed into a limited number of steps. A critical yet underexplored question is: what is the optimal noise to inject at each intermediate diffusion step? In this paper, we establish a theoretical framework that derives the closed-form analytical solution for optimal intermediate noise in diffusion models from a maximum likelihood estimation perspective, revealing a consistent conditional dependence structure that generalizes across diffusion paradigms. We instantiate this framework under the residual-shifting diffusion paradigm and accordingly design an LR-guided multi-input-aware noise predictor to replace random Gaussian noise. We further mitigate initialization bias with a high-quality pre-upsampling network. The compact 4-step trajectory uniquely enables end-to-end optimization of the entire reverse chain, which is computationally prohibitive for conventional long-trajectory diffusion models. Extensive experiments demonstrate that LPNSR achieves state-of-the-art perceptual performance on both synthetic and real-world datasets, without relying on any large-scale text-to-image priors. The source code of our method can be found at https://github.com/Faze-Hsw/LPNSR.
@@ 1 Introduction
We instantiate this framework on the residual-shifting diffusion paradigm, as its compact 4-step trajectory uniquely enables end-to-end optimization of the full reverse chain-computationally infeasible for long-trajectory models like DDPM (Ho et al., 2020). This allows us to learn a deep neural network for the theoretically optimal noise without modifying the pretrained denoising network or breaking the original efficient residual-shifting mechanism. We further address initialization bias from bicubic upsampling, a critical bottleneck for few-step sampling where limited iterations cannot correct initial deviations, via a pretrained pre-upsampling network to generate a high-quality initial state before diffusion. An additional key benefit of this design is that it enables arbitrary-step super-resolution inference without redesigning the diffusion framework’s hyper parameters or retraining the pretrained denoising network.
@@ 2 Related work
Image Super-Resolution. Along with the proliferation of deep learning, deep learning-driven approaches have progressively emerged as the dominant paradigm for SR (Dong et al., 2015; Rojas Sedó, 2022). Early prominent works primarily focused on training regression models using paired LR-HR data (Ahn et al., 2018; Kim et al., 2016; Wang et al., 2015). Though these models effectively capture the expectation of the posterior distribution, they inherently suffer from over-smoothing artifacts in generated results (Ledig et al., 2017; Menon et al., 2020; Sajjadi et al., 2017). To enhance the perceptual quality of reconstructed HR images, generative SR models have garnered growing interest, including autoregressive architectures (Dahl et al., 2017; Menick and Kalchbrenner, 2018; Van den Oord et al., 2016; Parmar et al., 2018). Despite notable gains in perceptual performance, autoregressive models typically incur substantial computational overhead. Additionally, GAN-based SR methods have attained remarkable success in perceptual quality (Guo et al., 2022; Karras et al., 2017; Ledig et al., 2017; Menon et al., 2020; Sajjadi et al., 2017), yet the training process remains unstable. More recently, diffusion-based models have become a focal point of SR research (Choi et al., 2021; Chung et al., 2022c; Kawar et al., 2022; Rombach et al., 2022; Saharia et al., 2023). These methods generally fall into two categories: those that concatenate the LR image to the denoiser’s input (Rombach et al., 2022; Saharia et al., 2023), and those that adapt the backward process of a pre-trained diffusion model (Choi et al., 2021; Chung et al., 2022c; Kawar et al., 2022). While these diffusion-based approaches yield promising performance, their methods still introduce unconstrained random Gaussian noise in each step of the reverse diffusion process, rather than meaningful noise maps.
@@
Diffusion Inversion. This paradigm aims to find the optimal noise maps that reconstruct the target image when fed into a diffusion model. Early works optimized text embeddings for better alignment (Gal et al., 2022; Mokady et al., 2023), and follow-up works further refined inversion via textual/visual prompts (Miyake et al., 2025; Nguyen et al., 2023) or intermediate noise map optimization (Ju et al., 2023; Kang et al., 2024; Meiri et al., 2023; Wallace et al., 2023). However, existing diffusion inversion methods are mostly heuristic step-wise optimization schemes tailored for image editing tasks. None of them establish a unified, generalizable theoretical derivation paradigm for optimal intermediate noise, nor provide a closed-form analytical solution that can generalize across mainstream diffusion paradigms. For SR tasks, InvSR (Yue et al., 2025) extends diffusion inversion to SR, but it is constrained by the long sampling trajectory of DDPM (Ho et al., 2020), which makes full-chain end-to-end optimization computationally prohibitive. It only optimizes the initial noise map, and fails to solve the fundamental problem of defining and optimizing the optimal intermediate noise for the full reverse sampling chain.
@@ 3.2 Theoretical derivation of optimal intermediate noise
Following standard diffusion posterior derivation practices (Ma et al., 2023), the parameterized distribution $p_{\theta}\left(x_{0}|y\right)$ learned by a well-trained diffusion model is a fitting to the data probability distribution $q\left(x_{0}|y\right)$. Thus, we adopt a non-informative prior $p_{\theta}\left(x_{0}|y\right)\propto 1$, and the marginal likelihood $p_{\theta}\left(x_{t-1}|y\right)$ is a normalization constant independent of $x_{0}$. This simplifies the posterior to $p_{\theta}\left(x_{0}|x_{t-1},y\right)\propto q\left(x_{t-1}|x_{0},y\right)$. Substituting the forward marginal Gaussian distribution in Eq. (2) and taking the logarithm, the log-likelihood becomes
$\log p_{\theta}\left(x_{0}\mid x_{t-1},y\right)\propto-\frac{\left\|x_{t-1}-\boldsymbol{\mu}_{t-1}^{m}(x_{0},y)\right\|^{2}}{2\Sigma_{t-1}^{m}}.$ (7)
Maximizing the log-likelihood is equivalent to minimizing the above L2 norm, leading to the optimization objective:
$z_{t-1}^{*}=\arg\min_{z_{t-1}}\left\|x_{t-1}(z_{t-1})-\boldsymbol{\mu}_{t-1}^{m}(x_{0},y)\right\|^{2}.$ (8)
@@ 3.3 Instantiation to residual-shifting diffusion SR paradigm
We instantiate our general optimal noise derivation paradigm proposed in Sec. 3.2 to the residual-shifting diffusion framework Yue et al. (2023), an efficient SR pipeline with a compact 4-step sampling trajectory. Unlike long-trajectory models (e.g., DDPM Ho et al. (2020)) requiring hundreds of steps, this 4-step design uniquely enables end-to-end optimization of the full reverse chain, which is computationally prohibitive for conventional long-trajectory diffusion models.
@@ 4.1 Experimental setup
Training Details. We train the noise predictor on the LSDIR Li et al. (2023) dataset and the first 10k face images from the FFHQ Karras et al. (2019) dataset for over 200k iterations, randomly cropping an image patch with a resolution of $256\times 256$ from the source image and synthesizing the LR image using the pipeline of RealESRGAN Wang et al. (2021) at each iteration. We adopt the AdamW Loshchilov and Hutter (2017) optimizer with a learning rate of $5\times 10^{-5}$ and a batch size of 16, while using the CosineAnnealing Loshchilov and Hutter (2016) scheduler with a minimum learning rate of $1\times 10^{-5}$. The hyperparameters for the loss function are set as $\lambda_{1}=1.0$, $\lambda_{l}=1.0$, and $\lambda_{g}=0.1$. During training, we set $T=4$ to remain consistent with ResShift Yue et al. (2023),and the noise variance hyperparameter $\kappa=2.0$ as well as the shifting sequence $\{\eta_{t}\}_{t=1}^{T}$ also follow the identical settings. The denoising network $f_{\theta}$ and the VQGAN autoencoder Esser et al. (2021) are frozen during training, only the noise predictor is optimized.
@@ 4.2 Experimental results
Inference Steps. Tab. 1 compares our PreSet-A (noise predictor only) and PreSet-B (with pre-upsampling) methods against original ResShift Yue et al. (2023) across 1-4 sampling steps. We observe that the pre-upsampling module delivers more significant gains with fewer steps, while our noise predictor stably improves perceptual performance across all settings, with all methods performing better with more steps. Regarding inference latency, our noise predictor introduces no noticeable overhead as its computation is performed entirely in the latent space. For 3-4 step inference, the model has sufficient iterations to mitigate initialization bias, so the pre-upsampling module can be optionally excluded to trade off a minor performance drop for faster inference. For 1-2 step inference, however, the pre-upsampling module is indispensable, as the limited sampling steps cannot compensate for initialization deviation without high-quality pre-upsampling. Qualitative results in Fig. 1 confirm that PreSet-B maintains consistent visual quality even in 1-2 step inference, while PreSet-A suffers from blurriness in low-step settings. We fix 4-step inference with pre-upsampling as our default LPNSR configuration for all following experiments.
@@
Performance Comparison. Tab. 2 and 3 present a comprehensive comparison of our LPNSR against recent SOTA methods on the ImageNet-Test, RealSR and RealSet80 datasets. Compared to the baseline ResShift Yue et al. (2023), our LPNSR achieves remarkable improvements in perceptual metrics (e.g., NIQE, CLIPIQA, MUSIQ) while maintaining competitive fidelity. Against T2I-utilizing models such as OSEDiff Wu et al. (2024a), InvSR Yue et al. (2025) and SeeSR Wu et al. (2024b), LPNSR delivers comparable or better perceptual quality without leveraging any pre-trained text-to-image priors. It also outperforms multi-step diffusion methods (e.g., SeeSR Wu et al. (2024b)) on core perceptual metrics. On real-world datasets, LPNSR ranks among the top-tier SOTA methods. It achieves leading perception-oriented metrics, such as NIQE, PI, CLIPIQA and MUSIQ on RealSR. On RealSet80, LPNSR attains top-2 PI score, the best MUSIQ and CLIPIQA scores among all competing methods. Qualitatively, Fig. 4 shows LPNSR generates sharper textures and more consistent structures than other methods, free from spurious details or over-smoothing (see Appendix for more visual comparisons). LPNSR generates SR images with sharp details, intact structural consistency, and no noticeable artifacts. It effectively restores natural textures and clear edge contours that align with the input LR structure, delivering visually coherent and realistic results.
@@
Analysis and Ablation Study of Noise Predictor. Across the 4-step coarse-to-fine denoising trajectory, our LR-guided noise predictor implements progressive prior guidance aligned with the denoising logic. As shown in Fig. 2, the predicted noise maps are highly aligned with the LR image’s structure and texture, following a hierarchical guidance pattern: the step 4 map anchors global structure to avoid initial sampling deviation, the step 3 map focuses on mid-frequency texture refinement to suppress cumulative error, and the step 2 map targets local fine-grained details to optimize perceptual quality. Statistical distribution analysis in Fig. 3 further confirms that the optimized noise distributions at all steps are not unconstrained random Gaussian distributions. We further conduct a step-wise ablation study on the RealSR dataset (see Tab. 4) to quantitatively verify the independent contribution of each step’s predictor: the full LPNSR model achieves the best overall performance, validating the effectiveness of our full-stage prior guidance; removing the step 4 predictor degrades both fidelity and perceptual quality, disabling the step 3 predictor causes the most severe PSNR drop, and replacing the step 2 predictor with random noise sharply impairs perceptual metrics, fully consistent with our qualitative and statistical observations.